For a new Kingsman movie which is going to be released soon, a producer of stylish watches has received the exclusive rights to sell a Kingsman watch The sales price, as negotiated in the contract, is $300, and a watch costs the producer $120 to assemble and ship to the stores. Their contract stipulates that 6 months after the Blu-Ray release of the movie, the unsold watches have to be discounted to the secondary market with an estimated price of $40. Assume that the producer is facing a one- shot production cycle and that it is impossible to produce additional watches later on. The demand for these watches is normally distributed with a mean of 2,000 and a standard deviation of 500. Determine the optimal number of watches to produce O About 2,750 watches O About 3,450 watches O About 1,850 watches About 2,250 watches
Answer #1

P=180/260=0.6923

z=0.5024

Q=mean+z*std dev

=2000+0.5024*500=2251.2  

about 2250 watches